OpenStudy (anonymous):

Multiply: x^2-x-42/x-8 * x^2-64/9x+54 Will medal!

OpenStudy (anonymous):

use the equation form to show us the equation properly. thx

OpenStudy (anonymous):

Then the answer is (x - 7) (x + 8)/ 9 . Am I correct?

OpenStudy (sachintha):

\[\frac{x^2-x-42}{x-8}\times\frac{x^2-64}{9x+54}\]

OpenStudy (sachintha):

You are correct. :D